OpenStudy (anonymous):

Simplify- X^-1 (X^7/-8) (X^1/8) Assume that all variables are positive

OpenStudy (tkhunny):

I wish I knew what Shay17 was doing. Rather just dropping answers (some correct) all over the place. @Shay17 please read the Code of Conduct and heed its wisdom. \(x^{-1}x^{-7/8}x^{1/8} = x^{-1 - (7/8) + (1/8)} = x^{-(8/8) - (7/8) + (1/8)} = x^{(-8-7+1)/8}\)

OpenStudy (tkhunny):

@Shay17 Please don't ignore the Code of Conduct, Someone needs to get your attention so you'll change your behavior.

OpenStudy (anonymous):

I don't need you telling me what to do @tkhunny and you still keep tagging me :/ stop

OpenStudy (tkhunny):

The Code of Conduct SHOULD be telling you what to do. Have you read it?

OpenStudy (anonymous):

hop off

OpenStudy (tkhunny):

I'll take that as a "no". Good day.

OpenStudy (anonymous):

\(\Large x^{-1} \times x^{\frac{7}{-8}} \times x^{\frac{1}{8}}\) =\(\Large x^{-1-\frac{7}{8} +\frac{1}{8}} = x^{\frac{-7}{4}}\)